OpenStudy (lucaz):

check my answer for this infinite geometric series: 1/3 - 1/9 + 1/27 - 1/81 ... I got 1/4

OpenStudy (anonymous):

20/81

OpenStudy (lucaz):

how did you get this?

OpenStudy (anonymous):

1/3 = 3/9 3/9 - 1/9 = 2/9 2/9 = 6/27 6/27 + 1/27 = 7/27 7/27 = 21/81 21/81 - 1/81 = 20/81 20/81

OpenStudy (lucaz):

it's an infinite series

OpenStudy (raden):

use this formula : S = a/(1 - r) here a = 1/3, r = -1/3

hartnn (hartnn):

yes, 1/3 / (1- (-1/3)) = 1/4 is correct

OpenStudy (lucaz):

ok, I did used that formula, wanted to ckeck.. thanks

OpenStudy (raden):

or if you want playing with algebra, it would be like this : let p = 1/3 - 1/9 + 1/27 - 1/81 + ... so, p = 1/3 - 1/3(1/3 - 1/9 + 1/27 + ...) p = 1/3 - 1/3p p + 1/3 p = 1/3 4/3 p = 1/3 p = 1/4 welcome :)

OpenStudy (lucaz):

cool, I didn't know this method, thank you!

OpenStudy (raden):

you're welcome

hartnn (hartnn):

thats also how that a/(1-r) standard formula was derived :)

OpenStudy (lucaz):

I've seen a different derivation, now I prefer this one